# Improving probabilistic error cancellation in the presence of non-stationary noise

### v1

@@ V-C Non-stationary PEC outcomes
The impact of the non-stationary noise can be seen in Fig. 8. The first plot, labeled No mitigation, presents the accuracy and stability metrics defined in Eqns. 16 and 19 respectively, for the raw Bernstein Vazirani circuit without any form of quantum error mitigation. The second plot, labeled ROEM, which stands for readout error mitigation, displays the metrics after performing SPAM noise mitigation. In this case, the SPAM noise parameters are held constant after initial device characterization. It deploys the standard matrix inversion [44] technique for mitigation. The third plot, labeled non-adaptive PEC, exhibits the accuracy and stability metrics for the Bernstein Vazirani circuit with non-adaptive PEC. This method incorporates SPAM noise mitigation within the PEC framework, as detailed in Sec. IV-C. The fourth plot, labeled adaptive PEC, presents the metrics for the adaptive PEC method, as discussed in Sec. III.
@@ VI Conclusion
However, scalability becomes an issue with the Bayesian approach in the presence of correlations. While the method can accurately estimate numerous noise parameters in the absence of correlations, scalability diminishes when dealing with highly correlated parameters due to the need to estimate joint distributions and perform Monte Carlo simulations for maximum-a-posterior optimization. In contemporary quantum devices, obtaining accurate noise characterization is difficult due to the non-stationary nature of noise and we emphasize the importance of integrating adaptive processes into PEC to effectively tackle non-stationary noise, especially considering its lengthy completion time. Refinement in noise model selection, optimization algorithms, and data collection frequencies can yield further improvements in probabilistic error cancellation and is a rich space to explore further.

### v2

@@ V-C Non-stationary PEC outcomes
The impact of the non-stationary noise can be seen in Fig. 8. The first plot, labeled No mitigation, presents the accuracy and stability metrics defined in Eqns. 20 and 23 respectively, for the raw Bernstein Vazirani circuit without any form of quantum error mitigation. The second plot, labeled ROEM, which stands for readout error mitigation, displays the metrics after performing SPAM noise mitigation. In this case, the SPAM noise parameters are held constant after initial device characterization. It deploys the standard matrix inversion [45] technique for mitigation. The third plot, labeled non-adaptive PEC, exhibits the accuracy and stability metrics for the Bernstein Vazirani circuit with non-adaptive PEC. This method incorporates SPAM noise mitigation within the PEC framework, as detailed in Sec. IV-C. The fourth plot, labeled adaptive PEC, presents the metrics for the adaptive PEC method, as discussed in Sec. III.
@@ VI Conclusion
Our goal in this paper was not to provide examples of performing Bayesian inference in different quantum computing contexts. Instead, we aimed to highlight that non-stationary noise is a significant issue in contemporary quantum computers, necessitating adaptive methods. Bayesian inference represents one such method. While our framework is generic, the specific equations for inference will differ based on the application. We illustrated our methodology using Bernstein-Vazirani as a test case for improving probabilistic error cancellation in presence of non-stationary noise.
Before concluding, we will discuss the scalability of the proposed algorithm. There are three potential areas where scalability may be a concern: PEC resource costs, optimization costs, and sampling costs.
A) PEC resource costs: The first scalability concern pertains to the number of PEC circuits required and is not directly related to Bayesian inference. The problem is well-known: each CNOT gate requires 16 PEC circuits for mitigating CNOT noise, and each qubit needs 2 PEC circuits for SPAM noise mitigation. Consequently, a Bernstein-Vazirani circuit with $m$ CNOT gates and $n$ qubits will need $2^{4m+n}$ PEC circuits. To mitigate this, one strategy is to avoid using PEC for SPAM mitigation and instead use standard matrix inversion (which may compromise some accuracy). Additionally, noise can be modeled in modular circuit layers, with PEC applied to these modules (instead of individual quantum gates).
B) Optimization: The second concern involves solving the non-linear system of equations (cf. Eqns. LABEL:eq:tv-correlated-means). In the Bernstein-Vazirani circuit, with $n$ qubits and $m$ CNOT gates, there are $2m+n$ noise parameters: each qubit contributes one SPAM parameter, and each CNOT gate’s asymmetric depolarizing noise model contributes $2m$ parameters. For an $n$-qubit problem, there can be $2^{n}$ such equations, one for each computational basis state. When $2m+n<2^{n}$, we have an over-determined system that can be solved using mean square error minimization. However, this is computationally expensive due to the high-dimensional search space for correlated parameters. If the system is under-determined, the estimates are unreliable due to multiple possible solutions. To make the algorithm scalable, focus on the most significant sources of noise and avoid overly detailed models that do not contribute significantly to noise explanation. For example, in this paper, we ignored gate noise in single-qubit gates and focused on SPAM noise and noise in entangling gates. In fact, even with quantum error correction, statistical mitigation methods may be relevant due to unforeseen noise sources. A single logical qubit may consist of thousands of physical qubits, so it is crucial to model noise at an aggregate, modular level (e.g., the logical level) rather than at lower levels of the computing stack. Choosing the right level of abstraction and avoiding overly granular noise models reduces the number of noise parameters that need to be estimated.
C) Sampling: Parameter estimation using Bayesian inference often requires sampling from a multivariate correlated distribution. However, the algorithm presented in this paper for estimating time-varying circuit noise parameters eliminates the need for sampling. As demonstrated in Eqn. 10, the parameters of the Dirichlet distribution, which models the joint density for the probability simplex, are updated using observed data $\mathcal{D}$, which help us formulate the system of non-linear equations to estimate the time-varying noise parameters. However, if custom metrics such as 95% confidence intervals are needed, or if we prefer using the maximum-a-posteriori (MAP) estimate instead of the mean, then Monte Carlo sampling will be required, which has a high computational overhead.